# Efficient Learning for Linear Properties of Bounded-Gate Quantum Circuits

### v1

@@ H.2 Proof of Corollary 1
Supported by the proposed ML model, the optimization of $\bm{\theta}$ can be entirely carried out on classical processors. More precisely, define $W(\bm{\theta})=U(\bm{x})$ (see Fig. J.8 for the visual interpretation), the two terms of the derivatives in Eq. (147) can be predicted by $h_{\mathsf{s}}$ in Alg. 1 and the optimized parameters can be obtained in an offline manner. For convenience, in the subsequent context, we dub the VQEs optimized by the proposed ML model as Offline-VQE. An alternative scenario involves utilizing the optimized parameters of Offline-VQE as effective initial parameters for running VQEs on real quantum devices. In either case, the offline optimization enabled by the proposed ML model dramatically reduces the demand for quantum resources, a critical advantage given the scarcity of available quantum processors in the current landscape.
Figure J.8: A toy model of optimizing variational quantum algorithms in an offline manner. Left panel: In the pre-training state, the learner collects training data from a $3$-qubit device with a hardware-efficient ansatz (HEA) $U(\bm{x})$, which contains $d=6$ tunable parameters. Then the learner can use the collected data to train an ML model introduced in the main text to enhance various downstream tasks. Middle panel: The optimization of variational quantum Eigen-solvers (VQEs) associated with the ansatz $W(\bm{\theta})$ can be accomplished by the proposed ML model without access to the quantum resources. Right panel: The optimization of quantum neural networks (QNNs) associated with the encoding circuit $V(\bm{z})$ and the trainable circuit $W(\bm{\theta})$ can be accomplished by the proposed ML model without the access to the quantum resources.
@@
where $\mathfrak{E}(g)$ is an optional regularizer. Given an unseen example $\bm{z}^{\prime}$, its predicted label is 00 if $g(\bm{z}^{\prime})<0.5$; otherwise, its predicted label is $1$.
@@
$\mathcal{G}=\left\{g(\bm{z},\bm{\theta})=\operatorname{Tr}\left(V(\bm{z})(\ket{0}\bra{0})^{\otimes N}V(\bm{z})^{\dagger}W(\bm{\theta})^{\dagger}\Pi_{0}W(\bm{\theta})\right)\Big{|}\bm{\theta}\in\Theta\right\},$ (149)
@@
Supported by the proposed ML model, the optimization of $\bm{\theta}$ for BQC can be entirely carried out on classical processors. According to Alg. 1, the classical input $\bm{x}$ should be divided into parts, where the first part refers to the training example $\bm{z}$ and the second part refers to the trainable parameters $\bm{\theta}$ in BiQCs. In addition, the circuit architecture of $U(\bm{x})$ amounts to the combination of $W(\bm{\theta})$ and $V(\bm{z})$ (see Fig. J.8 for the visual interpretation). In this way, for an arbitrary input $\bm{x}$ and trainable parameters $\bm{\theta}$, the derivatives $\partial\mathcal{L}/\partial\bm{\theta}_{k}$, which is formed by $g(\bm{z},\bm{\theta}^{(t)})$ and $g(\bm{z},\bm{\theta}^{(t,\pm)})$, can be estimated by the proposed ML model without accessing quantum resources. Due to this offline property, we call BiQCs optimized by the proposed ML model as Offline-BiQCs.
Remark. The approach of leveraging the proposed ML model to propel BiQCs can be readily extended to more complex scenarios. Specifically, the proposed ML model can enhance the optimization of multi-class quantum classifiers [121] and quantum regression models [123], significantly reducing the quantum resource overhead. Moreover, it can also improve quantum neural networks with diverse architectures, such as data re-uploading strategy [124] and quantum convolutional neural networks [125, 126].
Numerical results. To exhibit the effectiveness of our proposal for enhancing VQEs and QNNs, we employ it to estimate the ground state of the Transverse-field Ising (TFI) Hamiltonian and to classify a synthetic binary dataset, respectively. For both tasks, we consider the same quantum devices comprising $N=3$ qubits with a linear qubit connectivity, and the identical ansatz, i.e.,
@@
where $\operatorname{CNOT}_{a,b}$ denotes applying $\operatorname{CNOT}$ gate to $a$-th and $b$-th qubits. This ansatz contains in total $d=9$ classical inputs.
@@
Downstream task I: ground state energy estimation of TFI model. The mathematical form of the exploited $3$-qubit one-dimensional (1D) TFI Hamiltonian is $H_{\text{TFI}}=-0.1(Z_{1}Z_{2}+Z_{2}Z_{3})+0.5(X_{1}+X_{2}+X_{3})$. The ground state energy of $H_{\text{TFI}}$ is $\mathfrak{E}^{*}_{\text{TFI}}=-1.51\text{Ha}$.
Figure J.9: Simulation results for ground state energy estimation of TFI Hamiltonian. (a) The estimated energy of Offline-VQE under different hyper-parameter settings. The label $n=a$ refers that the number of training data is $a$. The label ‘Exact’ denotes the accurate ground state energy of the employed TFI Hamiltonian. The blue dotted line highlights the estimated result of traditional VQE with infinite shots. The vertical bar reflects the standard error of Offline-VQE in each setting. (b) The optimization process of Top-1 Offline-VQE and conventional VQE.
In numerical simulations, we adopt different hyper-parameter settings to evaluate the performance of Offline-VQE. In particular, we vary the number of training examples as $n\in\{30000,50000,70000,90000\}$ and set the truncation frequency as $\Lambda\in\{3,5,7,9\}$. The maximum iteration is set as $200$. The initial parameters are uniformly sampled from the range $[-\pi,\pi]^{9}$. Each setting is repeated $5$ times to obtain the statistical results. For comprehensive, we employ the conventional VQE with the same ansatz, the same initial parameters, the same optimizer, and infinite shots as the benchmark.
@@
Downstream task II: binary synthetic data classification. The construction rule of the binary synthetic dataset follows the study [118]. That is, we randomly and uniformly sample classical data $\bm{z}$ in the range $[-\pi,\pi]^{3}$ and then embed them into a $3$-qubit quantum circuit $V(\bm{z})=\operatorname{CNOT}_{2,3}\operatorname{CNOT}_{1,2}\bigotimes_{j=1}^{3}\operatorname{RY}(\bm{z}_{j})$. Then, we evolve the quantum state $V(\bm{z})\ket{0}^{\otimes 3}$ by the unitary $W(\bm{\theta}^{*})=\prod_{l=1}^{2}\left(\operatorname{CNOT}_{2,3}\operatorname{CNOT}_{1,2}\bigotimes_{j=1}^{3}\operatorname{RY}(\bm{\theta}^{*}_{j+3(l-1)})\right)$ with the fixed parameters $\bm{\theta}^{*}$ and measuring the evolved state under the observable $X\otimes\mathbb{I}_{4}$. For the input $\bm{z}^{(i)}$, its label is defined as
$y^{(i)}=\text{sign}\left(\operatorname{Tr}\left(W(\bm{\theta}^{*})V(\bm{z}^{(i)})(\ket{0}\bra{0})^{\otimes 3}V(\bm{z}^{(i)})^{\dagger}W(\bm{\theta}^{*})^{\dagger}\left(X\otimes\mathbb{I}_{4}\right)\right)\right),$ (151)
@@
In numerical simulations, we adopt different hyper-parameter settings to evaluate the performance of Offline-BiQC. The varied settings are analogous to those employed in the offline-VQE, where the number of training examples is set as $n\in\{30000,50000,70000\}$ and the truncation frequency is set as $\Lambda\in\{3,5,7,9\}$. The initial parameters are uniformly sampled from the range $[-\pi,\pi]^{6}$. The learning rate and the maximum iteration are set as $0.5$ and $140$, respectively. For comprehensive, we employ the conventional BiQC with the same encoding and training circuits, the same initial parameters, the same optimizer, and infinite shots as the benchmark.
@@
Figure J.10: Simulation results of Offline-BiQC. The figure depicts the test accuracy of Offline-BiQC under varied settings. The label ‘$\Lambda=a$&$n=b$’ denotes that the truncation value is $a$ and the number of training examples used to form the classical representations is $b$. The label ‘BiQC’ refers to the conventional BiQC optimized in an online manner.

### v2

@@ H.2 Proof of Corollary 1
Supported by the proposed ML model, the optimization of $\bm{\theta}$ can be entirely carried out on classical processors. More precisely, define $W(\bm{\theta})=U(\bm{x})$ (see Fig. 8 for the visual interpretation), the two terms of the derivatives in Eq. (151) can be predicted by $h_{\mathsf{s}}$ in Alg. 1 and the optimized parameters can be obtained in an offline manner. For convenience, in the subsequent context, we dub the VQEs optimized by the proposed ML model $h(\bm{x})$ as Offline-VQE. An alternative scenario involves utilizing the optimized parameters of Offline-VQE as effective initial parameters for running VQEs on real quantum devices. In either case, the offline optimization enabled by the proposed ML model dramatically reduces the demand for quantum resources, a critical advantage given the scarcity of available quantum processors in the current landscape.
Supplementary Fig. 8: A toy model of optimizing variational quantum algorithms in an offline manner. Left panel: In the pre-training state, the learner collects training data from a $3$-qubit device with a hardware-efficient ansatz (HEA) $U(\bm{x})$, which contains $d=6$ tunable parameters. Then the learner can use the collected data to train an ML model introduced in the main text to enhance various downstream tasks. Middle panel: The optimization of variational quantum Eigen-solvers (VQEs) associated with the ansatz $W(\bm{\theta})$ can be accomplished by the proposed ML model without access to the quantum resources. Right panel: The optimization of quantum neural networks (QNNs) associated with the encoding circuit $V(\bm{z})$ and the trainable circuit $W(\bm{\theta})$ can be accomplished by the proposed ML model without the access to the quantum resources.
@@
where $\mathfrak{E}(g)$ is an optional regularizer. Given an unseen example $\bm{z}^{\prime}$, its predicted label is 0 if $g(\bm{z}^{\prime})<0.5$; otherwise, its predicted label is $1$.
@@
[…]
@@
Supported by the proposed ML model, the optimization of $\bm{\theta}$ for BiQC can be entirely carried out on classical processors. According to Alg. 1, the classical input $\bm{x}$ should be divided into parts, where the first part refers to the training example $\bm{z}$ and the second part refers to the trainable parameters $\bm{\theta}$ in BiQCs. In addition, the circuit architecture of $U(\bm{x})$ amounts to the combination of $W(\bm{\theta})$ and $V(\bm{z})$ (see Fig. 8 for the visual interpretation). In this way, for an arbitrary input $\bm{z}$ and trainable parameters $\bm{\theta}$, the derivatives $\partial\mathcal{L}/\partial\bm{\theta}_{k}$, which is formed by $g(\bm{z},\bm{\theta}^{(t)})$ and $g(\bm{z},\bm{\theta}^{(t,\pm)})$, can be estimated by the proposed ML model without accessing quantum resources. Due to this offline property, we call BiQCs optimized by the proposed ML model as Offline-BiQCs.
Remark. The approach of leveraging the proposed ML model to propel BiQCs can be readily extended to more complex scenarios. Specifically, the proposed ML model can enhance the optimization of multi-class quantum classifiers [126] and quantum regression models [144], significantly reducing the quantum resource overhead. Moreover, it can also improve QNNs with diverse architectures, such as data re-uploading strategy [145] and quantum convolutional neural networks [146, 147].
K.4.1 Pretraining a 3-qubit HEA for both VQE and QNN
Recall that contemporary quantum devices encounter limitations such as connectivity, gate fidelities, and coherence time. To overcome these constraints, experimentalists typically adopt hardware-efficient ansatzes (HEAs) constructed from the native gate set from the specified quantum device to implement various variational quantum algorithms [132, 133]. Essentially, these HEAs are tailored to individual quantum devices, maintaining consistent gate layouts while adjusting parameters for various computational tasks, as delineated by the framework formulated in Eq. (3) of the main text. This intrinsic relation warrants the use of the proposed ML model to enhance plenty of variational quantum algorithms associated with device-specific HEAs.
To exhibit the effectiveness of our proposal for enhancing VQEs and QNNs, we employ it to pretrain a 3-qubit hardware-efficient ansatzes (HEAs) [132, 133], followed by estimating the ground state of the Transverse-field Ising model (TFIM) and classifying a synthetic binary dataset, respectively. The explicit form of the adopted HEA is
@@
where $\operatorname{CNOT}_{a,b}$ denotes applying $\operatorname{CNOT}$ gate to $a$-th and $b$-th qubits. This ansatz contains in total $d=9$ classical parameters.
@@
Downstream task I: ground state energy estimation of TFIM. The mathematical form of the exploited $3$-qubit one-dimensional (1D) TFIM is $\mathsf{H}_{\text{TFIM}}=-0.1(Z_{1}Z_{2}+Z_{2}Z_{3})+0.5(X_{1}+X_{2}+X_{3})$. The ground state energy of $\mathsf{H}_{\text{TFIM}}$ is $\mathfrak{E}^{*}_{\text{TFIM}}=-1.51\text{Ha}$.
Supplementary Fig. 9: Simulation results for ground state energy estimation of TFIM. (a) The estimated energy of Offline-VQE under different hyper-parameter settings. The label $n=a$ refers that the number of training data is $a$. The label ‘Exact’ denotes the accurate ground state energy of the employed TFIM. The blue dotted line highlights the estimated result of traditional VQE with infinite shots. The vertical bar reflects the standard error of Offline-VQE in each setting. (b) The optimization process of Top-1 Offline-VQE and conventional VQE.
In numerical simulations, we adopt different hyper-parameter settings to evaluate the performance of Offline-VQE. In particular, we vary the number of training examples as $n\in\{30000,50000,70000,90000\}$ and set the truncation frequency as $\Lambda\in\{3,5,7,9\}$. The maximum iteration is set as $200$. The initial parameters are uniformly sampled from the range $\pi,\pi]^{9}$. Each setting is repeated $5$ times to obtain the statistical results. For comprehensive, we employ the conventional VQE with the same ansatz, the same initial parameters, the same optimizer, and infinite shots as the benchmark.
The achieved results are depicted in Fig. 9. Fig. 9(a) demonstrates the performance of Offline-VQE under various settings. A key observation is that with the increased frequency truncation and number of training examples, the estimated ground state energy of Offline-VQE converges to the exact result. Namely, when $\Lambda\geq 7$ and $n\geq 70000$, the estimation error is less than $0.1\text{Ha}$. Besides, Fig. 9(b) compares the top-1 Offline-VQE (i.e., $\Lambda=7$ and $n=90000$) with conventional VQE. The achieved results exhibit a similar convergence rate during the optimization, where both of them converge to the near-optimal value after $50$ iterations. These results validate the potential of Offline-VQE in advancing conventional VQEs.
@@
[…]
@@
In numerical simulations, we adopt different hyper-parameter settings to evaluate the performance of Offline-BiQC. The varied settings are analogous to those employed in the offline-VQE, where the number of training examples is set as $n\in\{30000,50000,70000\}$ and the truncation frequency is set as $\Lambda\in\{3,5,7,9\}$. The initial parameters are uniformly sampled from the range $\pi,\pi]^{6}$. The learning rate and the maximum iteration are set as $0.5$ and $140$, respectively. For comprehensive, we employ the conventional BiQC with the same encoding and training circuits, the same initial parameters, the same optimizer, and infinite shots as the benchmark.
@@
Supplementary Fig. 10: Simulation results of Offline-BiQC. The figure depicts the test accuracy of Offline-BiQC under varied settings. The label $\Lambda=a$&$n=b$’ denotes that the truncation value is $a$ and the number of training examples used to form the classical representations is $b$. The label ‘BiQC’ refers to the conventional BiQC optimized in an online manner.